OpenStudy (anonymous):

Convert x = 4 to polar form.

OpenStudy (anonymous):

@Curry @ganeshie8

zepdrix (zepdrix):

Eyyy! So recall that when converting from cartesian to polar we use these, ya?\[\Large\rm x=r \cos \theta, \qquad\qquad y=r \sin \theta\]

zepdrix (zepdrix):

So the first step is pretty straight forward, ya?\[\Large\rm x=4\]\[\Large\rm r \cos \theta=4\]

zepdrix (zepdrix):

If we're able to, we would like to write this in a way that shows that r `is a function of` theta. So what can we do to `solve for r`. What do you think Mikeyyyyy? :)